# On the Invariant and Geometric Structure of the Holomorphic Unified Field Theory

### v1

@@ 5 Field Geometry and Global Spatial Topology
with a single HH–connection $\mathcal{A}=(\omega,A)$ whose curvature $\mathcal{F}=d\mathcal{A}+\mathcal{A}\!\wedge\!\mathcal{A}=(R,F)$ splits into spacetime curvature $R$ and Yang–Mills curvature $F$. On the complexified tangent bundle, we package the metric as a Hermitian field $g=h+iB$. On the real slice $y=0$ used for physics, the antisymmetric piece $B$ is not independent as it is identified with the gauge curvature two-form, $B\equiv F$, so the gauge-field shape is literally the two-form curvature living in the internal fiber.
@@
whose maximal torus or Cartan is $T^{4}=\mathrm{U}(1)^{4}$. The electric and hypercharge assignments live along these four commuting directions, the weights of the Cartan.
Spontaneous symmetry breaking proceeds in two steps:
@@
These spaces encode which gauge directions are unbroken versus broken, massless vs massive respectively.
Throughout this paper, we take spatial slices to be:
@@
Other constant-curvature topologies remain compatible with the framework such as a 3-torus $T^{3}$, periodic identifications on three cycles lead to discrete comoving momenta and possible global Aharonov–Bohm phases for gauge holonomies around non-contractible loops. Defect classification uses the same homotopy groups, but spatial infinity is replaced by large embedded 2-tori, 2-spheres, or by cycle representatives. A 3-sphere $S^{3}$, finite spatial volume, harmonic analysis uses $S^{3}$ eigenmodes. Large-radius $S^{2}\subset S^{3}$ still serves to define monopole charges; instanton number is unchanged since it lives in $\pi_{3}(G)$. But in all cases, the internal geometry, fiber $G$, Cartan $T^{4}$, vacuum manifolds, and the identification $B\equiv F$ on the real slice are unchanged, as only global boundary conditions and spectral discreteness differ.
@@
So in HUFT the universe’s field shape in this framework is a single connection on $P_{\mathrm{Spin}}\times_{M}P_{G}$ whose curvature splits into gravity and gauge parts, an internal $\mathrm{SU}(5)$ fiber containing a Cartan $T^{4}$ that organizes charges, vacuum manifolds $\mathcal{M}_{\mathrm{GUT}}$; flag space and $\mathcal{M}_{\mathrm{EW}}\cong S^{3}$ that determine massless vs. massive directions, topological sectors classified by maps from $S^{2}$ and $S^{3}$ into these spaces, and UV-smooth field textures controlled by $M_{*}$. For definiteness, we fix $\Sigma_{3}\simeq\mathbb{R}^{3}$, while noting that $T^{3}$ or $S^{3}$ spatial topologies are also compatible and leave the unification structure intact.
In the present formulation, HUFT requires that the real slice $M$ be an oriented, time-orientable spin 4–manifold supporting the product principal bundle $P_{\mathrm{tot}}=P_{\mathrm{Spin}}\times_{M}P_{G}$ with a single $H=\mathrm{Spin}(1,3)\times G$ connection $\mathcal{A}=(\omega,A)$ and curvature $\mathcal{F}=R\oplus F$. On the real slice, the antisymmetric metric piece satisfies $B\equiv F$, so the gauge sector is encoded directly in the fiber curvature. Chern–Weil matching and anomaly cancellation fix the normalization of $\langle\!\langle F\wedge F\rangle\!\rangle$ and select the minimal internal group $G=\mathrm{SU}(5)$. These are global bundle constraints on $M$, but they do not determine the global topology of spacelike slices $\Sigma_{t}$.
Accordingly, we do not commit the theory to a unique $\Sigma_{t}$ but common choices such as $\mathbb{R}^{3}$, $T^{3}$, or $S^{3}$ are all compatible, provided they admit the required spin structure and principal $G$-bundle. Physical consequences of different $\Sigma_{t}$ arise via boundary conditions such as discrete momenta on $T^{3}$, finite-volume harmonics on $S^{3}$, and possible global holonomies, not from the unification mechanism itself. Any further selection among allowed topologies would have to come from cosmological initial data or an extension of the action that weights topology classes.
@@ 6 Quantum Mechanics and Quantum Field Theory from HUFT
Now, to show compatibility with quantum mechanics, we will show the Born rule, Veltman condition, and Dirac equation come out naturally from fibre bundles in HUFT [36, 37, 38]. First, recall that $(M,g_{\mu\nu})$ is the complex Hermitian spacetime of HUFT with real slice $M\subset M_{\mathbb{C}}$. Matter fields live in a complex Hermitian vector bundle:
@@
and a unitary connection $D=\mathrm{d}+{\cal A}$ compatible with $h$. A pure quantum state at $x\in M$ is a ray $\psi_{x}]\in\mathbb{P}(E_{x})$, such as a nonzero vector modulo the $U(1)$ fiber phase. A ray $\psi]\in\mathbb{P}(E_{x})$ is the $1$-dimensional subspace $\{\lambda\psi:\lambda\in\mathbb{C}^{\times}\}$; physical predictions depend only on $\psi]$, equivalently on $\Pi_{[\psi]}=|\psi\rangle\langle\psi|/\langle\psi,\psi\rangle$. Observable $A$ is a Hermitian bundle endomorphism $A\in\Gamma(\mathrm{End}(E))$ at $x$ it has the spectral resolution $A_{x}=\sum_{i}a_{i}\,\Pi_{i,x}$ with orthogonal projectors $\Pi_{i,x}\in\mathrm{End}(E_{x})$.
Our goal is to construct a probability assignment with $\mu_{x}:\{\text{projectors on }E_{x}\}\to[0,1]$ meaning assign a probability number to every projector or outcome at $x$, such that for the outcome subspace $\mathrm{Im}\,\Pi_{i,x}$ we have:
@@
When $\dim E_{x}=2$, we replace A3 by $\sigma$-additivity for measurable fields of positive effects summing to $\mathbf{1}$ positive operator–valued measures (POVMs). By the Busch–Gleason extension [37], we again obtain: $\mu(E\,|\,\psi)=\mathrm{Tr}(\rho_{[\psi]}E)$ for all effects $E$, hence the same Born rule for projectors.
The preceding is fiberwise. To obtain spacetime probabilities for configuration-space localized outcomes, we must fix a Cauchy hypersurface $\Sigma\subset M$ with induced positive measure $\mathrm{d}\Sigma\sqrt{\gamma}$ from the real-slice metric. For a Hermitian line subbundle $L\subset E$ such as a position or detector mode like a localized wave-packet subbundle defined by a smooth section basis, the probability density is the fiber norm induced by $h$:
@@
Gauge covariance under the $U(1)$ factor of the HUFT structure group is guaranteed because $h$ and $D$ are unitary, $h(\psi,\psi)$ is $U(1)$-invariant, and parallel transport preserves the norm. Here $h(\psi,\psi)\equiv h_{x}(\psi_{x},\psi_{x})$ denotes the Hermitian fiber norm squared of $\psi$ at $x$, in a local frame $(h_{i\bar{j}})$ with $\psi=\psi^{i}e_{i}$ we have $h(\psi,\psi)=\bar{\psi}^{\bar{j}}h_{i\bar{j}}\psi^{i}$. Thus, the only $U(1)$-invariant quadratic functional compatible is the squared $h$-norm.
We let $E_{x}=E^{S}_{x}\otimes E^{E}_{x}$ describe system and environment fibers. For a Schmidt state $|\psi\rangle=\sum_{k}\sqrt{p_{k}}\,|k\rangle_{S}\otimes|k\rangle_{E}$, HUFT’s unitary fiber symmetries contain phase twirls $U_{S}\otimes U_{E}^{\dagger}$ that leave $|\psi\rangle$ invariant called environment-assisted invariance [38]. Equal-weight cases, $p_{k}$ equal force equiprobability on the system outcomes by symmetry, rational weights follow by refinement, and continuity from (A3) yields $\mu(\Pi_{k}\,|\,\psi)=p_{k}=\|\Pi_{k}\psi\|^{2}/\|\psi\|^{2}.$ Within HUFT, states are rays in a Hermitian vector bundle, and measurements are Hermitian bundle endomorphisms. Assuming ray invariance, additivity on orthogonal outcomes, and locality or noncontextuality for subspaces, the unique probability assignment is:
@@
The geometric set-up as before holds, but we as well assume a ${\rm Spin}^{c}$ structure to derive the Dirac equation. We let $S\to M$ be the complex spinor bundle with Hermitian fiber metric $h_{S}$ and Clifford map $c:T^{*}M\to{\rm End}(S)$, $c(\alpha)c(\beta)+c(\beta)c(\alpha)=2\,h^{\mu\nu}\alpha_{\mu}\beta_{\nu}\mathbf{1}$, meaning $c$ is a map $c:T^{\ast}M\to\mathrm{End}(S)$. For any one-forms $\alpha,\beta$:
@@
Equivalently, with $c(dx^{\mu})=\gamma^{\mu}$, this reads $\{\gamma^{\mu},\gamma^{\nu}\}=2\,h^{\mu\nu}\mathbf{1}$, as this is the gamma-matrix anticommutation relation. For the connection we let $\nabla^{\rm LC}$ be the Levi–Civita connection of $h_{\mu\nu}$, $\omega^{ab}{}_{\mu}$ its spin connection, and $A_{\mu}\equiv A_{\mu}^{A}T_{A}+A^{(1)}_{\mu}$ the $G\times U(1)$ gauge potential. The total unitary covariant derivative on sections of $S\otimes E_{\rm rep}$ is:
@@
Compatibility means $D_{\mu}h_{S}=0$ and $D_{\mu},c(e_{\nu})]=\Gamma^{\rho}_{\mu\nu}c(e_{\rho})$, this is equivalent to metric or Clifford compatibility and ensures $D_{\mu}h_{S}=0$.. We consider the Dirac Lagrangian density on the real slice [3, 29]:
@@
for $\psi\in\Gamma(S\otimes E_{\rm rep})$, this read as
$\psi$ is a section of the tensor–product bundle $S\otimes E_{\rm rep}$, meaning at each spacetime point at $x\in M$:
@@
so $\psi$ is a spinor field from $S$ that as well carries an internal gauge index from $E_{\rm rep}$. Here $\Gamma(\cdot)$ denotes the space of smooth sections. In components, one can think of $\psi^{ai}(x)$ with spinor index $a$ and internal representation index $i$. Varying with respect to $\bar{\psi}$ and using metric or Clifford compatibility gives:
@@
This means the Dirac equation is precisely the covariant, unitary, Clifford-compatible section equation on the spinor bundle determined by HUFT’s geometry and gauge structure. But on the complex slice, we would use the holomorphic or anti-holomorphic split, but restriction to $M$ yields the above.
@@
where $\mathcal{E}\in\Gamma({\rm End}(\mathcal{V}))$ is an endomorphism built from background fields such as Higgs potential curvature, Yukawa endomorphisms, covariant curvatures of $D_{\mu}$, and $\mathcal{V}$ is the vector bundle that the fluctuation field lives in. The regulated one-loop effective action can be written via the heat kernel as [13, 14]:
@@
a supertrace over all fluctuating species, bosons with $(+)$ sign, fermions with ()(-) sign, including ghosts.
Specializing to backgrounds where only $H$ is nonzero and slowly varying, $\mathcal{E}$ restricted to each species reduces at leading order to a mass-squared endomorphism that is affine in $H^{\dagger}H$:
@@
with $c_{\text{sp}}$ a group-theory or counting factor and $\lambda_{\text{sp}}$ the relevant coupling. The coefficient of the induced operator $H^{\dagger}H$ in the quadratically divergent part of $\Gamma^{(1)}$ is proportional to the fiberwise supertrace:
@@
where the displayed terms are, respectively, the Higgs self-coupling, ${\rm SU}(2)$ and ${\rm U}(1)_{Y}$ gauge couplings, and the top Yukawa, the ellipsis denotes smaller Yukawas and any additional fields in the chosen fiber content.
The Veltman condition is treated as the fiber-geometric statement that this supertrace vanishes at some renormalization scale $\mu$:
@@
The scale $\mu$ where ${\rm STr}\,\partial_{H^{\dagger}H}\mathcal{E}(\mu)=0$ is scheme-dependent so the supertrace structure itself is the universal $a_{1}$ coefficient. The interpretation of this is $a_{1}\propto{\rm STr}\,\mathcal{E}$ is a bundle trace of an endomorphism, so the sum rule is just the statement that the quadratic counterterm to $H^{\dagger}H$ disappears when the supertrace over all associated bundles, with correct statistics and ghost structure, is zero. In HUFT with entire-function regulators, quadratic divergences are tamed, but the same $a_{1}$ coefficient governs the finite threshold correction replacing the naive $\Lambda^{2}$ piece, so the supertrace relation remains the geometric criterion for suppressing the Higgs mass renormalization.
From the same geometric structure as before, we define a single holomorphic principal bundle [44]:
@@
where $\omega$ is the complex spin connection, $R$ the complexified Riemann curvature, and $A,F$ the internal gauge connection and curvature. We let $e$ denote the complex soldering form; on the real slice $e$ induces $h=g_{(\mu\nu)}$.
We now choose an $\mathrm{Ad}$-invariant, nondegenerate bilinear form on the Lie algebra:
@@
compatible with anomaly cancellation and boundary conditions specified below. With this normalization, we define the one-coupling holomorphic action:
$S_{\mathrm{HUFT}}=\Re\int_{M}\Bigg[\frac{1}{g_{\star}^{2}}\,\langle\mathcal{F},\star\mathcal{F}\rangle_{\mathcal{H}}\ +\ \lambda\,\langle B,\mathcal{F}\rangle_{\mathcal{H}}\ +\ \mu\,\langle e\wedge e,\ R\rangle_{\mathrm{Lor}}\ +\ S_{\mathrm{matter}}(\mathcal{A},e;\,\Psi)\Bigg],$ (100)
where $B$ is an auxiliary holomorphic $\mathfrak{h}$-valued two-form and $\star$ is the complex Hodge operator. Varying (100) with respect to $B$ gives:
@@
where the last statement follows after imposing reality or Hodge conditions on $(M,h)$. Eliminating $B$ yields:
$S_{\mathrm{HUFT}}=\Re\int_{M}\Big[\frac{1}{g_{\star}^{2}}\,\langle\mathcal{F},\star\mathcal{F}\rangle_{\mathcal{H}}+\mu\,\langle e\wedge e,\ R\rangle_{\mathrm{Lor}}+S_{\mathrm{matter}}\Big].$ (102)
@@
The entire-function regulator $f(\Box/M_{*}^{2})$ preserves the holomorphic Ward and Slavnov–Taylor identities [8, 9, 10], yielding a common renormalization above $M_{*}$:
@@
To show the internal gauge symmetries from our fibre bundle we let $E\to M$ be a rank-$n$ holomorphic Hermitian vector bundle for matter, with $c_{1}(E)=0$, a nowhere-vanishing holomorphic volume form $\Omega_{E}\in H^{0}(M,\wedge^{n}E^{*})$, and assume $E$ is slope-stable so that the compatible connection is Hermitian–Yang–Mills. Preservation of $(h_{E},\Omega_{E})$ reduces the internal structure group to $\mathrm{SU}(n)$, and the internal connection has holonomy contained in $\mathrm{SU}(n)$.
We demand a single adjoint reduction to the real-slice group, correct hypercharge quantization obtained from an integral pairing derived from (99), and chiral, anomaly-free matter for one family from associated bundles of $E$. Then the minimal rank is $n=5$.
Now for the mixed Ward identities to show the Diff $\leftrightarrow$ Gauge interlock we fix holomorphic gauge conditions for $\mathcal{A}$ and diffeomorphisms on $M$. To preserve the gauge conditions, an infinitesimal diffeomorphism $\delta_{\xi}$ must be accompanied by a compensating internal transformation $\delta_{\epsilon}$:

### v2

@@ 5 Field Geometry and Global Spatial Topology
with a single $H$-connection $\mathcal{A}=(\omega,A)$ whose curvature $\mathcal{F}=d\mathcal{A}+\mathcal{A}\!\wedge\!\mathcal{A}=(R,F)$ splits into spacetime curvature $R$ and Yang–Mills curvature $F$. On the complexified tangent bundle, we package the metric as a Hermitian field $g=h+iB$. On the real slice $y=0$ used for physics, the antisymmetric piece $B$ is not independent, as it is identified with the gauge curvature two-form, $B\equiv F$, so the gauge-field shape is literally the two-form curvature living in the internal fiber.
@@
whose maximal torus or Cartan is $T^{4}=\mathrm{U}(1)^{4}$. The electric and hypercharge assignments live along these four commuting directions, the weights of the Cartan. Spontaneous symmetry breaking proceeds in two steps:
@@
These spaces encode which gauge directions are unbroken versus broken, massless vs, massive, respectively. Throughout this paper, we take spatial slices to be:
@@
Other constant-curvature topologies remain compatible with the framework, such as a three-torus $T^{3}$, where periodic identifications on three cycles lead to discrete comoving momenta and possible global Aharonov–Bohm phases for gauge holonomies around non-contractible loops. Defect classification uses the same homotopy groups, but spatial infinity is replaced by large embedded two-tori, two-spheres, or by cycle representatives. A three-sphere $S^{3}$, finite spatial volume, harmonic analysis uses $S^{3}$ eigenmodes. Large-radius $S^{2}\subset S^{3}$ still serves to define monopole charges; instanton number is unchanged since it lives in $\pi_{3}(G)$. But in all cases, the internal geometry, fiber $G$, Cartan $T^{4}$, vacuum manifolds, and the identification $B\equiv F$ on the real slice are unchanged, as only global boundary conditions and spectral discreteness differ.
@@
So, in HUFT, the universe’s field shape in this framework is a single connection on $P_{\mathrm{Spin}}\times_{M}P_{G}$ whose curvature splits into gravity and gauge parts, an internal $\mathrm{SU}(5)$ fiber containing a Cartan $T^{4}$ that organizes charges, vacuum manifolds $\mathcal{M}_{\mathrm{GUT}}$; flag space and $\mathcal{M}_{\mathrm{EW}}\cong S^{3}$ that determine massless vs. massive directions, topological sectors classified by maps from $S^{2}$ and $S^{3}$ into these spaces, and UV-smooth field textures controlled by $M_{*}$. For definiteness, we fix $\Sigma_{3}\simeq\mathbb{R}^{3}$, while noting that $T^{3}$ or $S^{3}$ spatial topologies are also compatible and leave the unification structure intact.
In the present formulation, HUFT requires that the real slice $M$ be an oriented, time-orientable spin four-manifold supporting the product principal bundle $P_{\mathrm{tot}}=P_{\mathrm{Spin}}\times_{M}P_{G}$ with a single $H=\mathrm{Spin}(1,3)\times G$ connection $\mathcal{A}=(\omega,A)$ and curvature $\mathcal{F}=R\oplus F$. On the real slice, the antisymmetric metric piece satisfies $B\equiv F$, so the gauge sector is encoded directly in the fiber curvature. Chern–Weil matching and anomaly cancellation fix the normalization of $\langle\!\langle F\wedge F\rangle\!\rangle$ and select the minimal internal group $G=\mathrm{SU}(5)$. These are global bundle constraints on $M$, but they do not determine the global topology of spacelike slices $\Sigma_{t}$.
Accordingly, we do not commit the theory to a unique $\Sigma_{t}$, but common choices such as $\mathbb{R}^{3}$, $T^{3}$, or $S^{3}$ are all compatible, provided they admit the required spin structure and principal $G$-bundle. Physical consequences of different $\Sigma_{t}$ arise via boundary conditions such as discrete momenta on $T^{3}$, finite-volume harmonics on $S^{3}$, and possible global holonomies, not from the unification mechanism itself. Any further selection among allowed topologies would have to come from cosmological initial data or an extension of the action that weights topology classes.
@@ 6 Quantum Mechanics and Quantum Field Theory from HUFT
Now, to show compatibility with quantum mechanics, we will show that the Born rule, Veltman condition, and Dirac equation come out naturally from fibre bundles in HUFT Gleason1957 ; BuschLahtiBook ; ZurekEnvariance ; Veltman1981 . First, recall that $(M,g_{\mu\nu})$ is the complex Hermitian spacetime of HUFT with real slice $M\subset M_{\mathbb{C}}$. Matter fields live in a complex Hermitian vector bundle:
@@
and a unitary connection $D=\mathrm{d}+{\cal A}$ compatible with $h$. A pure quantum state at $x\in M$ is a ray $\psi_{x}]\in\mathbb{P}(E_{x})$, such as a nonzero vector modulo the $U(1)$ fiber phase. A ray $\psi]\in\mathbb{P}(E_{x})$ is the $1$-dimensional subspace $\{\lambda\psi:\lambda\in\mathbb{C}^{\times}\}$; physical predictions depend only on $\psi]$, equivalently on $\Pi_{[\psi]}=|\psi\rangle\langle\psi|/\langle\psi,\psi\rangle$. Observable $A$ is a Hermitian bundle endomorphism $A\in\Gamma(\mathrm{End}(E))$; at $x$, it has the spectral resolution $A_{x}=\sum_{i}a_{i}\,\Pi_{i,x}$ with orthogonal projectors $\Pi_{i,x}\in\mathrm{End}(E_{x})$.
Our goal is to construct a probability assignment with $\mu_{x}:\{\text{projectors on }E_{x}\}\to[0,1]$ meaning that we assign a probability number to every projector or outcome at $x$, such that for the outcome subspace $\mathrm{Im}\,\Pi_{i,x}$, we have
@@
When $\dim E_{x}=2$, we replace A3 by $\sigma$-additivity for measurable fields of positive effects summing to $\mathbf{1}$ positive operator-valued measures (POVMs). By the Busch–Gleason extension BuschLahtiBook , we again obtain $\mu(E\,|\,\psi)=\mathrm{Tr}(\rho_{[\psi]}E)$ for all effects $E$; hence, we have the same Born rule for projectors.
The preceding information is fiberwise. To obtain spacetime probabilities for configuration-space localized outcomes, we must fix a Cauchy hypersurface $\Sigma\subset M$ with induced positive measure $\mathrm{d}\Sigma\sqrt{\gamma}$ from the real-slice metric. For a Hermitian line subbundle $L\subset E$ such as a position or detector mode like a localized wave-packet subbundle defined by a smooth section basis, the probability density is the fiber norm induced by $h$:
@@
Gauge covariance under the $U(1)$ factor of the HUFT structure group is guaranteed because $h$ and $D$ are unitary, $h(\psi,\psi)$ is $U(1)$-invariant, and parallel transport preserves the norm. Here, $h(\psi,\psi)\equiv h_{x}(\psi_{x},\psi_{x})$ denotes the Hermitian fiber norm squared of $\psi$ at $x$; in a local frame $(h_{i\bar{j}})$ with $\psi=\psi^{i}e_{i}$, we have $h(\psi,\psi)=\bar{\psi}^{\bar{j}}h_{i\bar{j}}\psi^{i}$. Thus, the only $U(1)$-invariant quadratic functional compatible is the squared $h$-norm.
We let $E_{x}=E^{S}_{x}\otimes E^{E}_{x}$ describe system and environment fibers. For a Schmidt state $|\psi\rangle=\sum_{k}\sqrt{p_{k}}\,|k\rangle_{S}\otimes|k\rangle_{E}$, HUFT’s unitary fiber symmetries contain phase twirls $U_{S}\otimes U_{E}^{\dagger}$ that leave $|\psi\rangle$ invariant called environment-assisted invariance ZurekEnvariance . For equal-weight cases, $p_{k}$ equal force equiprobability on the system outcomes by symmetry, rational weights follow by refinement, and continuity from (A3) yields $\mu(\Pi_{k}\,|\,\psi)=p_{k}=\|\Pi_{k}\psi\|^{2}/\|\psi\|^{2}.$ Within HUFT, states are rays in a Hermitian vector bundle, and measurements are Hermitian bundle endomorphisms. Assuming ray invariance, additivity on orthogonal outcomes, and locality or noncontextuality for subspaces, the unique probability assignment is:
@@
The geometric setup described before holds, but we also assume a ${\rm Spin}^{c}$ structure to derive the Dirac equation. We let $S\to M$ be the complex spinor bundle with Hermitian fiber metric $h_{S}$ and Clifford map $c:T^{*}M\to{\rm End}(S)$, $c(\alpha)c(\beta)+c(\beta)c(\alpha)=2\,h^{\mu\nu}\alpha_{\mu}\beta_{\nu}\mathbf{1}$, meaning $c$ is a map $c:T^{\ast}M\to\mathrm{End}(S)$. For any one-forms $\alpha,\beta$:
@@
Equivalently, with $c(dx^{\mu})=\gamma^{\mu}$, this reads $\{\gamma^{\mu},\gamma^{\nu}\}=2\,h^{\mu\nu}\mathbf{1}$, as this is the gamma-matrix anticommutation relation. For the connection, we let $\nabla^{\rm LC}$ be the Levi–Civita connection of $h_{\mu\nu}$, $\omega^{ab}{}_{\mu}$ its spin connection, and $A_{\mu}\equiv A_{\mu}^{A}T_{A}+A^{(1)}_{\mu}$ the $G\times U(1)$ gauge potential. The total unitary covariant derivative on sections of $S\otimes E_{\rm rep}$ is:
@@
Compatibility means $D_{\mu}h_{S}=0$ and $D_{\mu},c(e_{\nu})]=\Gamma^{\rho}_{\mu\nu}c(e_{\rho})$; this is equivalent to metric or Clifford compatibility and ensures $D_{\mu}h_{S}=0$. We consider the Dirac Lagrangian density on the real slice WaldGR ; ParkerToms :
@@
for $\psi\in\Gamma(S\otimes E_{\rm rep})$, this reads as $\psi$ is a section of the tensor–product bundle $S\otimes E_{\rm rep}$, meaning at each spacetime point at $x\in M$:
@@
so $\psi$ is a spinor field from $S$ that also carries an internal gauge index from $E_{\rm rep}$. Here, $\Gamma(\cdot)$ denotes the space of the smooth sections. In components, one can think of $\psi^{ai}(x)$ with spinor index $a$ and internal representation index $i$. Varying with respect to $\bar{\psi}$ and using metric or Clifford compatibility gives:
@@
This means the Dirac equation is the precise covariant, unitary, Clifford-compatible section equation on the spinor bundle determined by HUFT’s geometry and gauge structure. On the complex slice, we would use the holomorphic or anti-holomorphic split, but restriction to $M$ yields the above.
@@
where $\mathcal{E}\in\Gamma({\rm End}(\mathcal{V}))$ is an endomorphism built from background fields such as Higgs potential curvature, Yukawa endomorphisms, covariant curvatures of $D_{\mu}$, and $\mathcal{V}$ is the vector bundle that the fluctuation field lives in. The regulated one-loop effective action can be written via the heat kernel as DeWitt1965 ; Vassilevich2003 :
@@
a supertrace over all fluctuating species, bosons with $(+)$ sign, and fermions with ()(-) sign, including ghosts. Specializing in backgrounds where only $H$ is nonzero and slowly varying, $\mathcal{E}$ restricted to each species reduces at leading order to a mass-squared endomorphism that is affine in $H^{\dagger}H$:
@@
with $c_{\text{sp}}$ a group-theory or counting factor and $\lambda_{\text{sp}}$ is the relevant coupling. The coefficient of the induced operator $H^{\dagger}H$ in the quadratically divergent part of $\Gamma^{(1)}$ is proportional to the fiberwise supertrace:
@@
where the displayed terms are, respectively, the Higgs self-coupling, ${\rm SU}(2)$ and ${\rm U}(1)_{Y}$ gauge couplings, and the top Yukawa; the ellipsis denotes smaller Yukawas and any additional fields in the chosen fiber content.
The Veltman condition is treated as the fiber-geometric statement that this supertrace vanishes at some renormalization scale $\mu$ Veltman1981 :
@@
The scale $\mu$ where ${\rm STr}\,\partial_{H^{\dagger}H}\mathcal{E}(\mu)=0$ is scheme-dependent so the supertrace structure itself is the universal $a_{1}$ coefficient. The interpretation of this is $a_{1}\propto{\rm STr}\,\mathcal{E}$ is a bundle trace of an endomorphism, so the sum rule is just the statement that the quadratic counterterm to $H^{\dagger}H$ disappears when the supertrace over all associated bundles, with correct statistics and ghost structure, is zero. In HUFT with entire-function regulators, quadratic divergences are tamed, but the same $a_{1}$ coefficient governs the finite threshold correction, replacing the naive $\Lambda^{2}$ piece, so the supertrace relation remains the geometric criterion for suppressing the Higgs mass renormalization.
Using the same geometric structure as before, we define a single holomorphic principal bundle MoffatThompsonEmbedding2025 :
@@
where $\omega$ is the complex spin connection, $R$ the complexified Riemann curvature, and $A,F$ the internal gauge connection and curvature. We let $e$ denote the complex soldering form; on the real slice, $e$ induces $h=g_{(\mu\nu)}$. We now choose an $\mathrm{Ad}$-invariant, nondegenerate bilinear form on the Lie algebra:
@@
compatible with the anomaly cancellation and boundary conditions specified below. With this normalization, we define the one-coupling holomorphic action:
$S_{\mathrm{HUFT}}=\mathrm{Re}\int_{M}\Bigg[\frac{1}{g_{\star}^{2}}\,\langle\mathcal{F},\star\mathcal{F}\rangle_{\mathcal{H}}\ +\ \lambda\,\langle B,\mathcal{F}\rangle_{\mathcal{H}}\ +\ \mu\,\langle e\wedge e,\ R\rangle_{\mathrm{Lor}}\ +\ S_{\mathrm{matter}}(\mathcal{A},e;\,\Psi)\Bigg],$ (120)
where $B$ is an auxiliary holomorphic $\mathfrak{h}$-valued two-form and $\star$ is the complex Hodge operator. Varying (120) with respect to $B$ gives
@@
where the last statement follows after imposing reality or Hodge conditions on $(M,h)$. Eliminating $B$ yields
$S_{\mathrm{HUFT}}=\mathrm{Re}\int_{M}\Big[\frac{1}{g_{\star}^{2}}\,\langle\mathcal{F},\star\mathcal{F}\rangle_{\mathcal{H}}+\mu\,\langle e\wedge e,\ R\rangle_{\mathrm{Lor}}+S_{\mathrm{matter}}\Big].$ (122)
@@
The entire-function regulator $f(\Box/M_{*}^{2})$ preserves the holomorphic Ward and Slavnov–Taylor identities Slavnov1972 ; Taylor1971 ; BRST1976 , yielding a common renormalization above $M_{*}$
@@
To show the internal gauge symmetries from our fibre bundle, we let $E\to M$ be a rank-$n$ holomorphic Hermitian vector bundle for matter, with $c_{1}(E)=0$, a nowhere-vanishing holomorphic volume form $\Omega_{E}\in H^{0}(M,\wedge^{n}E^{*})$, and assume $E$ is slope-stable so that the compatible connection is Hermitian–Yang–Mills
The dynamical HUFT action and field equations are formulated on the Lorentzian real slice $(M,h)$. The appeal to slope stability and Donaldson–Uhlenbeck–Yau is not an additional hypothesis on $(M,h)$; rather, it is a standard holomorphic-bundle criterion applied on an auxiliary compact Kähler base $(X,\omega)$ associated with the internal/matter bundle data. Restricting back to the real slice selects the same reduced unitary structure group such as $SU(n)$ when $c_{1}(E)=0$), while the Lorentzian dynamics proceed entirely with the usual gauge-natural constructions on $(M,h)$. Preservation of $(h_{E},\Omega_{E})$ reduces the internal structure group to $\mathrm{SU}(n)$, and the internal connection has holonomy contained in $\mathrm{SU}(n)$.
We demand a single adjoint reduction to the real-slice group, correct hypercharge quantization obtained from an integral pairing derived from (119), and chiral, anomaly-free matter for one family from associated bundles of $E$. Then, the minimal rank is $n=5$.
Now, for the mixed Ward identities to show the Diff $\leftrightarrow$ Gauge interlock, we fix holomorphic gauge conditions for $\mathcal{A}$ and diffeomorphisms on $M$. To preserve the gauge conditions, an infinitesimal diffeomorphism $\delta_{\xi}$ must be accompanied by a compensating internal transformation $\delta_{\epsilon}$: